John was born two years after his father's marriage.his mother is 5 years younger than his father but 20 years older than John who is 10 years old. how old was his father when he was married?
Answer by ankor@dixie-net.com(22740)

John was born two years after his father's marriage.
his mother is 5 years younger than his father but 20 years older than John who is 10 years old.
how old was his father when he was married?
:
From the given information," mother is 5 years younger than his father but 20 years older than John who is 10 years old.", we know:
Mother's age: 20 + 10 = 30 yrs old
and
Father's age: 30 + 5 = 35 yrs old:
:
"John was born two years after his father's marriage..." how old was his father when he was married?"
John was born 10 year ago, therefore father was married 12 yrs ago
35 - 12 = 23 yrs old when father was married